COMMUTATOR STRUCTURE OF OPERATOR IDEALS

KEN DYKEMA TADEUSZ FIGIEL GARY WEISS MARIUSZ WODZICKI

Abstract. We describe the commutator structure of ideals in the algebra of bounded operators on an infinitely-dimensional separable Hilbert space.